# Speech-FT: Merging Pre-trained And Fine-Tuned Speech Representation Models For Cross-Task Generalization

### v2

@@ III-C Evaluation
Datasets and tasks. To evaluate the cross-task generalization ability of speech representations, we use the SUPERB [6] benchmark, which assesses models across a diverse set of downstream tasks, covering content, speaker, paralinguistic, and semantic aspects. Evaluating the entire SUPERB benchmark is computationally expensive due to its wide range of tasks. MiniSUPERB [37] addresses this by selecting representative tasks, demonstrating that evaluation on a well-chosen subset sufficiently reflects a model’s performance. Following this approach, we evaluate four tasks: phoneme recognition (PR, content) on Librispeech [27], speaker identification (SID, speaker) on Voxceleb1 [31], emotion recognition (ER, paralinguistic) on IEMOCAP [32], and slot filling (SF, semantic) on Audio SNIP [38]. To provide a more comprehensive evaluation, we additionally assess full SUPERB benchmark performance on representative models in Section V-A. We report phone error rate (PER%) for PR, accuracy (ACC%) for SID and ER, and slot-type F1 score (F1) along with slot value CER (CER%) for SF. We adhere to the default SUPERB evaluation setup, where the speech representation model remains frozen, and the weighted sum of its layer-wise features is used as input for the downstream model. Recall from Section II-A that our goal is to evaluate cross-task generalization. As such, the task prediction model $D$ is discarded after fine-tuning. Instead, we train a separate downstream model for each SUPERB evaluation task. During training, only the learnable weighted sum and the downstream model are updated. To ensure the stability of the weighted sum features, we apply layer normalization to the representations before computing the sum. The performance of HuBERT, wav2vec 2.0, DeCoAR 2.0, and WavLM Base+ reported in this paper has been reproduced with this modification. We adhere to the default number of update steps specified by the official SUPERB evaluation tool, s3prl [^5] https://github.com/s3prl/s3prl/tree/main, for downstream model training.
@@
SUPERB Score. To facilitate a more intuitive and unified comparison of different speech representation models, we adopt the SUPERB score ($\text{SUPERB}_{\text{S}}$) as introduced in the SUPERB SLT 2022 Challenge [8]. This score linearly scales each task’s performance between a traditional baseline (FBank) and the best known result (SOTA) from prior work, reflecting how much each model improves upon traditional features relative to state-of-the-art representations. When traditional spectral features already perform close to SOTA, even small improvements are treated as more significant—implicitly accounting for task difficulty. For tasks with multiple evaluation metrics, an average is first computed across metrics within the same task, followed by an average across all tasks. Mathematically, let $\phi_{t,j}(f)$ be the score of model $f$ on metric $j$ of task $t$, with $\phi_{t,j}(\text{Baseline})$ and $\phi_{t,j}(\text{SOTA})$ denoting the corresponding baseline and SOTA values, respectively. Define $\mathcal{T}$ as the set of evaluation tasks and $\mathcal{M}_{t}$ as the set of metrics for task $t$. The full formulation is:
@@
The performance of both FBank and SOTA models is obtained from the SUPERB journal extension [39]. The final score is scaled by a factor of 1000 for readability. In this paper, we report $\text{SUPERB}_{\text{S}}$ based on four downstream tasks: PR, SID, ER, and SF. This metric has been widely adopted in previous studies for evaluating generalization across speech tasks [40, 41, 42, 43, 44, 45].
TABLE IV: Results of Speech-FT on the unsupervised fine-tuning task, where HuBERT is continuously pre-trained on the Chinese speech corpus AISHELL-3 [33].
Method English Chinese
PR SID ER SF $\text{SUPERB}_{\text{S}}\uparrow$ zh-CN ASR
PER%$\downarrow$ ACC%$\uparrow$ ACC%$\uparrow$ F1$\uparrow$ CER%$\downarrow$ CER%$\downarrow$
Speech-FT 5.27 81.57 65.43 87.82 25.20 866.51 24.23
Stable-FT 10.84 78.52 62.81 85.28 30.44 789.88 23.47
Pre-trained 5.17 81.86 64.99 88.54 24.70 870.2 24.94
TABLE V: Results of Speech-FT when using two fine-tuning tasks: phoneme recognition (PR) on Librispeech and speaker identification (SID) on VoxCeleb1. The background color highlights cases where the evaluation task is included in the fine-tuning tasks.
@@
TABLE VI: The results of Speech-FT under a gradually increasing number of fine-tuning tasks. “PR” denotes PR on Librispeech, “SID” denotes SID on Voxceleb1, and “ER” denotes ER on IEMOCAP.
@@
Baselines. Our goal is to improve pre-trained speech representation models through fine-tuning. Therefore, the most fundamental baseline is the pre-trained model itself, denoted as Pre-trained. Another key baseline is stable fine-tuning, abbreviated as Stable-FT. We exclude regular fine-tuning from our baselines for the following reasons: First, we empirically found that in some cases, regular fine-tuning fails entirely, preventing successful fine-tuning of the speech representation model. Second, even in cases where regular fine-tuning succeeds, it consistently underperforms compared to Stable-FT, making Stable-FT a stronger and more reliable baseline, as shown in Table I.
@@
In addition to these two core comparisons, we further benchmark Speech-FT against several fine-tuning strategies that constrain weight deviation during fine-tuning, as detailed in Section V-C. These strategies include: (1) Weight-space regularization, which penalizes deviations from the pre-trained weights during fine-tuning; (2) LoRA fine-tuning [17], which inserts trainable low-rank adapters instead of modifying the full set of parameters; (3) Early stopping, where checkpoints are taken early in the fine-tuning process to limit weight deviation.
@@ IV Main Results
TABLE VII: Results of Speech-FT on wav2vec 2.0, DeCoAR 2.0, and WavLM Base+ when fine-tuning with ASR on TED-LIUM.
Model Method PR SID ER SF $\text{SUPERB}_{\text{S}}\uparrow$
PER%$\downarrow$ ACC%$\uparrow$ ACC%$\uparrow$ F1$\uparrow$ CER%$\downarrow$
Wav2vec 2.0 Speech-FT 4.14 78.44 64.87 88.37 24.1 864.79
Stable-FT 6.84 72.95 64.53 87.22 26.45 824.82
Pre-trained 5.78 75.26 63.45 88.12 24.88 837.08
DeCoAR 2.0 Speech-FT 10.29 73.98 63.01 85.37 30.39 781.77
Stable-FT 16.92 66.56 61.16 83.14 34.47 701.6
Pre-trained 14.91 74.21 62.46 83.18 34.66 736.77
WavLM Base+ Speech-FT 3.42 88.12 69.02 90.73 21.1 945.5
Stable-FT 3.93 83.2 67.63 89.41 22.95 907.31
Pre-trained 3.94 89.43 68.67 90.52 21.3 942.96
TABLE VIII: The full SUPERB benchmark evaluation results for fine-tuning HuBERT and WavLM Base+ with ASR on TED-LIUM. Note that the SUPERB score “$\text{SUPERB}_{\text{full}}$” reported here is computed across the entire SUPERB benchmark and, therefore, is not directly comparable to the $\text{SUPERB}_{\text{s}}$ scores in this paper, which are based on only four evaluation tasks.
@@
WavLM Base+ Speech-FT 3.42 5.18 97.44 0.1090 88.12 4.34 3.84 69.02 99.10 90.73 21.10 950.6
@@ IV-A Supervised Fine-tuning
We begin by fine-tuning HuBERT on various supervised tasks, including automatic speech recognition (ASR) on TED-LIUM [24], phoneme classification (PC) on TIMIT [25], speaker identification (SID) on Librispeech [27], emotion recognition (ER) on CREMA-D [29], ASR on Librispeech, phoneme recognition (PR) on Librispeech, SID on VoxCeleb1 [31], and ER on IEMOCAP [32].
Table II presents the SUPERB score results, where Speech-FT consistently outperforms Stable-FT across all supervised fine-tuning tasks. Additionally, Speech-FT improves performance compared to the pre-trained model for all tasks.
A detailed breakdown of the SUPERB downstream results is provided in Table III. The background color highlights cases where the fine-tuning task matches the evaluation task. Overall, Stable-FT significantly degrades performance on unrelated tasks. For instance, fine-tuning on SID with VoxCeleb1 causes substantial degradation in PR and SF, with PER increasing by 23.91% and CER by 18.57%, respectively. Similarly, fine-tuning on SID with Librispeech leads to performance drops in PR and SF, with PER increasing by 9.01% and CER by 6.84%. Fine-tuning on PC with TIMIT negatively impacts SID, reducing ACC by 15.52%.
In contrast, Speech-FT not only preserves the cross-task generalization ability of the pre-trained model but also enhances overall representation quality, as reflected in the increased SUPERB score. Notably, fine-tuning on ASR tasks such as TED-LIUM yields consistent improvements across all evaluation tasks. We hypothesize that this is because ASR involves learning both phonetic and temporal patterns, which benefit other tasks through shared feature representations. For example, Speech-FT on TED-LIUM leads to a 1.23% PER reduction on PR, a 2.25% ACC increase on SID, a 2.79% ACC increase on ER, and a 1.24% CER reduction on SF. These results indicate that Speech-FT effectively leverages rich ASR supervision to improve speech representations.
@@ IV-B Unsupervised Fine-tuning
We further evaluate the effectiveness of Speech-FT in an unsupervised fine-tuning scenario, where no task-specific labels are available. Specifically, we simulate a cross-lingual setting by continuing HuBERT pre-training on AISHELL-3 [33], a large-scale Mandarin speech corpus, using the self-supervised objective described in Equation 7. This setting mimics the realistic case where practitioners adapt pre-trained speech models to new domains or languages using only unlabeled data.
To evaluate whether the model retains its ability to represent the original language, we assess performance on English downstream tasks from SUPERB. Additionally, to examine adaptation to the new language, we introduce Mandarin ASR from the Common Voice corpus [46] as an auxiliary evaluation task.
Results are shown in Table IV. We observe that continued pre-training alone (i.e., Stable-FT) severely degrades generalization to English tasks, causing a drop of 80.32 in the SUPERB score. In contrast, Speech-FT achieves a significantly better trade-off: it improves performance on the Chinese ASR task while largely preserving the model’s English ability. This demonstrates that Speech-FT mitigates the representational forgetting typically observed in naive unsupervised adaptation and enables more effective cross-lingual fine-tuning.
@@ IV-C Multiple Fine-tuning Tasks
The results are shown in Table V. Overall, Stable-FT significantly degrades performance on PR, ER, and SF. For example, under the sequential fine-tuning strategy, PR increases by 20.55% PER, ER decreases by 4.96% ACC, and SF increases by 18.06% CER. Similarly, with the TIES merging strategy, PR increases by 9.5% PER, ER decreases by 2.53% ACC, and SF increases by 7.75% CER. In contrast, Speech-FT effectively mitigates this degradation, leading to an overall improvement in the SUPERB score compared to the pre-trained model.
@@
To further illustrate the benefits of Speech-FT in multiple fine-tuning tasks, we present the results of gradually increasing the number of fine-tuning tasks in Table VI. In this experiment, we adopt linear merging as a representative strategy due to its practical advantage of efficiently utilizing already fine-tuned models without requiring joint retraining. The results indicate that as more fine-tuning tasks are introduced, Speech-FT steadily improves the SUPERB score, demonstrating its ability to absorb and leverage information from different tasks.
TABLE IX: Ablation studies on stable fine-tuning in Speech-FT. Background color highlights cases where the fine-tuning task matches the evaluation task. We use phoneme classification (PC) on TIMIT and speaker identification (SID) on Voxceleb1 as the fine-tuning tasks.
@@ V-A Generality of Speech-FT across Models and Evaluation Tasks
To investigate the generality of Speech-FT, we first examine Speech-FT across speech representation models with diverse pre-training objectives, and then assess its performance under a comprehensive evaluation setting using the full SUPERB benchmark [6]. We adopt ASR on TED-LIUM [24] as the fine-tuning task throughout this subsection.
Speech-FT is effective across different speech representation models.

### v3

@@ III-C Evaluation
Datasets and tasks. To evaluate the cross-task generalization ability of speech representations, we use the SUPERB [yang2021superb] benchmark, which assesses models across a diverse set of downstream tasks, covering content, speaker, paralinguistic, and semantic aspects. Evaluating the entire SUPERB benchmark is computationally expensive due to its wide range of tasks. MiniSUPERB [wang2023minisuperb] addresses this by selecting representative tasks, demonstrating that evaluation on a well-chosen subset sufficiently reflects a model’s performance. Following this approach, we evaluate four tasks: phoneme recognition (PR, content) on Librispeech [panayotov2015librispeech], speaker identification (SID, speaker) on Voxceleb1 [nagrani2020voxceleb], emotion recognition (ER, paralinguistic) on IEMOCAP [busso2008iemocap], and slot filling (SF, semantic) on Audio SNIP [lai2021semi]. To provide a more comprehensive evaluation, we additionally assess full SUPERB benchmark performance on representative models in Section V-A. We report phone error rate (PER%) for PR, accuracy (ACC%) for SID and ER, and slot-type F1 score (F1) along with slot value CER (CER%) for SF. We adhere to the default SUPERB evaluation setup, where the speech representation model remains frozen, and the weighted sum of its layer-wise features is used as input for the downstream model. Recall from Section II-A that our goal is to evaluate cross-task generalization. As such, the task prediction model $D$ is discarded after fine-tuning. Instead, we train a separate downstream model for each SUPERB evaluation task. During training, only the learnable weighted sum and the downstream model are updated. To ensure the stability of the weighted sum features, we apply layer normalization to the representations before computing the sum. The performance of HuBERT, wav2vec 2.0, DeCoAR 2.0, and WavLM Base+ reported in this paper has been reproduced with this modification. We adhere to the default number of update steps specified by the official SUPERB evaluation tool, s3prl [^4] https://github.com/s3prl/s3prl/tree/main, for downstream model training.
@@
SUPERB Score. To facilitate a more intuitive and unified comparison of different speech representation models, we adopt the SUPERB score ($\text{SUPERB}_{\text{S}}$) as introduced in the SUPERB SLT 2022 Challenge [feng2023superb]. This score linearly scales each task’s performance between a traditional baseline (FBank) and the best known result (SOTA) from prior work, reflecting how much each model improves upon traditional features relative to state-of-the-art representations. When traditional spectral features already perform close to SOTA, even small improvements are treated as more significant—implicitly accounting for task difficulty. For tasks with multiple evaluation metrics, an average is first computed across metrics within the same task, followed by an average across all tasks. Mathematically, let $\phi_{t,j}(f)$ be the score of model $f$ on metric $j$ of task $t$, with $\phi_{t,j}(\text{Baseline})$ and $\phi_{t,j}(\text{SOTA})$ denoting the corresponding baseline and SOTA values, respectively. Define $\mathcal{T}$ as the set of evaluation tasks and $\mathcal{M}_{t}$ as the set of metrics for task $t$. The full formulation is:
@@
The performance of both FBank and SOTA models is obtained from the SUPERB journal extension [yang2024large]. The final score is scaled by a factor of 1000 for readability. In this paper, we report $\text{SUPERB}_{\text{S}}$ based on four downstream tasks: PR, SID, ER, and SF. This metric has been widely adopted in previous studies for evaluating generalization across speech tasks [shi2023ml, shi2023multi, wang2023task, chang2024colld].
TABLE II: Results of Speech-FT on the unsupervised fine-tuning tasks, where we continue HuBERT pre-training on the Chinese speech corpus AISHELL-3 [shi2020aishell], the Spanish subset of the Multilingual LibriSpeech dataset [pratap2020mls], and the English dataset VoxCeleb1 [nagrani2020voxceleb]. Stable-FT (20%) denotes the snapshot of the Stable-FT model saved at 20% of the total fine-tuning steps.
Language Method English Chinese Spanish
zh-CN ASR es ASR
[…]
Chinese (cross-lingual) Speech-FT 866.51 24.23 -
Stable-FT 789.88 23.47 -
Stable-FT (20%) 832.37 24.26 -
Spanish (cross-lingual) Speech-FT 868.26 - 32.87
Stable-FT 747.50 - 30.36
English (mono-lingual) Speech-FT 876.10 - -
Stable-FT 867.96 - -
Pre-trained 870.20 24.94 35.39
TABLE III: Results of Speech-FT when using two fine-tuning tasks: phoneme recognition (PR) on Librispeech and speaker identification (SID) on VoxCeleb1. The background color highlights cases where the evaluation task is included in the fine-tuning tasks.
@@
TABLE IV: The results of Speech-FT under a gradually increasing number of fine-tuning tasks. “PR” denotes PR on Librispeech, “SID” denotes SID on Voxceleb1, and “ER” denotes ER on IEMOCAP.
@@
Baselines. Our goal is to improve pre-trained speech representation models through fine-tuning. Therefore, the most fundamental baseline is the pre-trained model itself, denoted as Pre-trained. Another key baseline is stable fine-tuning, abbreviated as Stable-FT. We exclude regular fine-tuning from our baselines for the following reasons: First, we empirically found that in some cases, regular fine-tuning fails entirely, preventing successful fine-tuning of the speech representation model. Second, even in cases where regular fine-tuning succeeds, it consistently underperforms compared to Stable-FT, making Stable-FT a stronger and more reliable baseline [^5] For example, on speaker identification (Voxceleb1), Stable-FT achieves a SUPERB score of 651.42 versus 568.62 with regular fine-tuning, and on emotion recognition (CREMA-D), Stable-FT achieves 847.33 versus 834.38..
@@
In addition to these two core comparisons, we further benchmark Speech-FT against several fine-tuning strategies that constrain weight deviation during fine-tuning, as detailed in Section V-C. These strategies include: (1) Weight-space regularization, which penalizes deviations from the pre-trained weights during fine-tuning; (2) Parameter-efficient fine-tuning [hu2022lora, liu2024dora], which inserts trainable lightweight adapters instead of modifying the full set of parameters; (3) Early stopping, where checkpoints are taken early in the fine-tuning process to limit weight deviation.
@@ IV Main Results
TABLE V: Results of Speech-FT on wav2vec 2.0, DeCoAR 2.0, and WavLM Base+ when fine-tuning with ASR on TED-LIUM.
Model Method $\text{SUPERB}_{\text{S}}\uparrow$
wav2vec 2.0 Speech-FT 864.79
Stable-FT 824.82
Pre-trained 837.08
DeCoAR 2.0 Speech-FT 781.77
Stable-FT 701.60
Pre-trained 736.77
WavLM Base+ Speech-FT 945.50
Stable-FT 907.31
Pre-trained 942.96
TABLE VI: The full SUPERB benchmark evaluation results for fine-tuning HuBERT and WavLM Base+ with ASR on TED-LIUM. Note that the SUPERB score $\text{SUPERB}_{\text{full}}$” reported here is computed across the entire SUPERB benchmark and, therefore, is not directly comparable to the $\text{SUPERB}_{\text{s}}$ scores in this paper, which are based on only four evaluation tasks.
@@
WavLM Base+ Speech-FT 3.42 5.18 97.44 0.1090 88.12 4.34 3.84 69.02 99.10 90.73 21.10 950.60
@@ IV-A Supervised Fine-tuning
We begin by fine-tuning HuBERT on various supervised tasks, including automatic speech recognition (ASR) on TED-LIUM [rousseau2012ted], phoneme classification (PC) on TIMIT [garofolo1993darpa], speaker identification (SID) on Librispeech [panayotov2015librispeech], emotion recognition (ER) on CREMA-D [cao2014crema], ASR on Librispeech, phoneme recognition (PR) on Librispeech, SID on VoxCeleb1 [nagrani2020voxceleb], and ER on IEMOCAP [busso2008iemocap].
Table I presents the SUPERB score results, where Speech-FT consistently outperforms Stable-FT across all supervised fine-tuning tasks. Additionally, Speech-FT improves performance compared to the pre-trained model for all tasks. The table further provides a detailed breakdown of the downstream results, with the background color highlighting cases in which the fine-tuning task matches the evaluation task. Overall, Stable-FT significantly degrades performance on unrelated tasks. For instance, fine-tuning on SID with VoxCeleb1 causes substantial degradation in PR and SF, with PER increasing by 23.91% and CER by 18.57%, respectively. Similarly, fine-tuning on SID with Librispeech leads to performance drops in PR and SF, with PER increasing by 9.01% and CER by 6.84%. Fine-tuning on PC with TIMIT negatively impacts SID, reducing ACC by 15.52%.
In contrast, Speech-FT not only preserves the cross-task generalization ability of the pre-trained model but also enhances overall representation quality, as reflected in the increased SUPERB score. Notably, fine-tuning on ASR tasks such as TED-LIUM yields consistent improvements across all evaluation tasks. We hypothesize that this is because ASR involves learning both phonetic and temporal patterns, which benefit other tasks through shared feature representations. For example, Speech-FT on TED-LIUM leads to a 1.23% PER reduction on PR, a 2.25% ACC increase on SID, a 2.79% ACC increase on ER, and a 1.24% CER reduction on SF. These results indicate that Speech-FT effectively leverages rich ASR supervision to improve speech representations. It is also worth noting that fine-tuning on IEMOCAP achieves the highest SUPERB score (1010.29), mainly due to the large improvement on ER (from 64.99% to 85.95%). For PR, SID, and SF, fine-tuning on ASR offers substantially greater benefits.
@@ IV-B Unsupervised Fine-tuning
We further evaluate the effectiveness of Speech-FT in an unsupervised fine-tuning scenario, where no task-specific labels are available. Specifically, we simulate a cross-lingual setting by continuing HuBERT pre-training on AISHELL-3 [shi2020aishell], a large-scale Mandarin speech corpus. This setting mimics the realistic case where practitioners adapt pre-trained speech models to new domains or languages using only unlabeled data.
To evaluate whether the model retains its ability to represent the original language, we assess performance on English downstream tasks from SUPERB. Additionally, to examine adaptation to the new language, we introduce Mandarin ASR from the Common Voice corpus [ardila2019common] as an auxiliary evaluation task.
Results are shown in Table II. We observe that continued pre-training alone (i.e., Stable-FT) severely degrades generalization to English tasks, causing a drop of 80.32 in the SUPERB score. In contrast, Speech-FT achieves a balanced trade-off: it improves performance on the Chinese ASR task while largely preserving the model’s English ability. Compared with an earlier Stable-FT checkpoint (20% of fine-tuning steps), Speech-FT yields similar Chinese ASR results but maintains substantially higher English performance, further highlighting its advantage. This demonstrates that Speech-FT mitigates the representational forgetting typically observed in naive unsupervised adaptation and enables more effective cross-lingual fine-tuning.
In addition to Chinese, we also extend the setting to Spanish by continuing HuBERT pre-training on the Spanish subset of the Multilingual LibriSpeech dataset [pratap2020mls] and evaluating Spanish ASR on the Common Voice corpus [ardila2019common]. For Spanish, Speech-FT works better than in the Chinese setting, likely because Spanish is closer to the pretrained language, which is English. Furthermore, we conduct mono-lingual unsupervised fine-tuning on VoxCeleb1, which contains in-the-wild interview speech with diverse speakers and noisy conditions. The results show that Speech-FT also improves the representation quality in this setting.
@@ IV-C Multiple Fine-tuning Tasks
The results are shown in Table III. Overall, Stable-FT significantly degrades performance on PR, ER, and SF. For example, under the sequential fine-tuning strategy, PR increases by 20.55% PER, ER decreases by 4.96% ACC, and SF increases by 18.06% CER. Similarly, with the TIES merging strategy, PR increases by 9.5% PER, ER decreases by 2.53% ACC, and SF increases by 7.75% CER. In contrast, Speech-FT effectively mitigates this degradation, leading to an overall improvement in the SUPERB score compared to the pre-trained model.
@@
To further illustrate the benefits of Speech-FT in multiple fine-tuning tasks, we present the results of gradually increasing the number of fine-tuning tasks in Table IV. In this experiment, we adopt linear merging as a representative strategy due to its practical advantage of efficiently utilizing already fine-tuned models without requiring joint retraining. The results indicate that as more fine-tuning tasks are introduced, Speech-FT steadily improves the SUPERB score, demonstrating its ability to absorb and leverage information from different tasks.
It is also important to note that multitask fine-tuning does not always outperform strong single-task baselines on every evaluation. For example, the PR+SID model yields slightly lower SID accuracy compared to the SID-only model, indicating that negative interference between tasks can indeed occur. Nevertheless, the PR+SID model achieves a higher overall SUPERB score (886.36) than the SID-only model (881.07) as shown in Table IV. These findings demonstrate that while task-level trade-offs exist, Speech-FT in multitask settings can still enhance overall representation quality.
TABLE VII: Ablation studies on stable fine-tuning in Speech-FT. Background color highlights cases where the fine-tuning task matches the evaluation task. We use phoneme classification (PC) on TIMIT and speaker identification (SID) on Voxceleb1 as the fine-tuning tasks.
@@ V-A Generality of Speech-FT across Models and Evaluation Tasks
To investigate the generality of Speech-FT, we first examine Speech-FT across speech representation models with diverse pre-training objectives, and then assess its performance under a comprehensive evaluation setting using the full SUPERB benchmark [yang2021superb]. We adopt ASR on TED-LIUM [rousseau2012ted] as the fine-tuning task throughout this subsection.
V-A1 Speech-FT is effective across different speech representation models